OpenStudy (anonymous):

Bernice is selling seashells she has found at the beach. The price of shell depends on the length of the shell. Cm 5 8 12 20 25 Price ($) 2 3.5 5 18 40 A.)find the exponential model for the data B.)what is the length of a shell selling for 9$ C.)40cm conch shell should sell for him much?

OpenStudy (mathmale):

It's important to know the basic form of the "exponential model." While y=e^x is the simplest form, in approaching a problem such as this one, I prefer y=a*e^(bx), where a and b are constants to be determined.

OpenStudy (anonymous):

So what would A and B represent?

OpenStudy (mathmale):

We can assume that this data does follow an exponential model: the problem statement asks us to find such a model. We have two unknowns, a and b, in this model. To find a and b, we need to develop two simultaneous equations in a and b. Pick any two x values in the given table, along with the corresponding y values. For example, pick (8,3.5) and (25,40). Write out y=a*e^(bx) in each case, substituting the relevant x for a and the relevant y for b. For example: y=a*e^(bx) becomes 3.5=a*e(8b). Solve the resulting system of two equations in the two unknowns, a and b. Substitution will probably prove to be the easiest way of doing this.

OpenStudy (anonymous):

Ok that makes sense,thanks for being patient!

OpenStudy (mathmale):

a represents the value y would have if x were = to zero: y=a*e^0. b in this case would be called the "exponential growth constant;" it has to do with how fast y increases as x increases.

OpenStudy (mathmale):

Happy to work with you!

OpenStudy (anonymous):

That's exactly what I'm learning,great!